# DECO: Life-Cycle Management of Enterprise-Grade Chatbots

### v1

@@ 6.2. IcM Retrieval Improvement
  • Search Type: A list specifying the scope of the search, which can include one or more of the following values:
@@
      By title: Searches the codebase by title, useful for locating where specific elements like methods or classes are defined.
@@
      By content: Searches the codebase by content, applicable when seeking definitions of specific methods or classes.
@@
      By reference: Searches the codebase by reference, targeting locations where specific methods or classes are referenced.
@@
6.3. TSG Retrieval Improvement
The get_tsg function is a critical component of our backend flow, supported by an efficient and accurate TSG retrieval algorithm. Traditional retrieval-augmented generation (RAG) approaches typically use similarity-based searches, which compare user query embeddings with document chunk embeddings or employ term-frequency techniques (Robertson et al., 2009). However, empirical analyses show that these methods often underperform with highly technical documents, as embeddings do not capture the subtle distinctions within technical terms effectively. Although recent advancements in supervised retrieval models, such as RAFT (Zhang et al., 2024b), AR2 (Zhang et al., 2022), and ARG (Sachan et al., 2023), have improved retrieval accuracy, they necessitate the development and fine-tuning of deep neural networks (DNNs). These approaches are impractical for our self-hosting framework in DECO, as they require specialized machine learning expertise and additional infrastructure for training and hosting ML models, which may not be readily available to tenant teams.
To address these challenges, we developed a hybrid search algorithm integrated with Azure AI Search, leveraging an advanced preprocessing pipeline detailed in Section 4, eliminating the need for additional ML training while enhancing both retrieval accuracy and computational efficiency. Key improvements to the TSG retrieval process are detailed below:
Query Rephrasing User queries are rephrased using the Semantic Kernel in the planner to optimize search text, generating search arguments. Searches are conducted using both the original user query and the generated search arguments.
Repository Description During the preprocessing pipeline setup, tenants provide brief descriptions for each Git repository or document source. The descriptions are included with the retrieved documents from the corresponding sources to enhance the LLM’s contextual understanding about the document.
Document Filtering An additional filtering step refines the search results by evaluating the similarity scores of the top documents. If the top document(s) demonstrates a significant margin in similarity scores over subsequent results, only the top document(s) are returned. This approach, based on the observation that a small set of highly relevant TSGs usually suffices to resolve issues, prevents the inclusion of overly detailed or irrelevant documents. Advanced filtering techniques are further discussed in (Anonymous, 2024b).
Multi-Field Hybrid Search The hybrid search algorithm utilizes the preprocessing pipeline to extract metadata from each document chunk, including raw content, title, and the associated published URL (as detailed in Section 4). These metadata elements serve as separate embedding targets, and the outputs from all search queries are combined to effectively filter document chunks. We evaluated various search strategies such as vector search, hybrid search, and semantic search, ultimately adopting a hybrid approach that combines embeddings from titles and raw content for optimal performance, based on reciprocal rank fusion (RRF) (Cormack et al., 2009; Microsoft Corporation, 2024). Future enhancements include incorporating advanced retrieval algorithms like HyDE and Reverse HyDE (Gao et al., 2022), generating enriched matching fields, and exploring the integration of graph structures into the index (Anonymous, 2024a) and leveraging feedback from historical conversations to dynamically refine search rankings (Anonymous, 2024b). Section 8.2 evaluates the efficiency and effectiveness of these improvements.
@@ 7. Frontend
The frontend of DECO is built around the Azure Bot, which leverages the Azure Bot Framework V4 and is deployed on an Azure Web Application (see Figure 7). This setup coordinates API calls, manages messaging (interacting with the PromptFlow endpoint), retains state in Azure storage, and supports user interface enhancements through frontend plugins such as Adaptive Cards (Microsoft, 2024a), allowing users to modify generated Kusto queries. Authentication is handled via Microsoft Entra ID, also known as Azure Active Directory (AAD) authentication (see Figure 7). Different groups, managed by Azure Core Identity, are used to restrict user access.
The web application connects users to the backend Azure Bot through direct line web channels, which communicate with the Azure Bot API. This web channel seamlessly integrates with the IcM portal, pre-populating inputs like ticket IDs for backend processing. Additionally, the bot can be accessed via Microsoft Teams through a published manifest. The web application also transmits system telemetry to a reporting dashboard for monitoring. Detailed interactions between the components are shown in Figure 7.
@@ 7.1. Plugin
We introduce frontend “Plugins” to support various types of user interfaces required by certain customized skills. For example, for the kql_generator skill, the backend output includes plugin data that triggers a frontend plugin, which displays an Adaptive Card for the user. This card allows users to select Kusto clusters from a dropdown menu, edit the query, use checkboxes to choose which queries to execute, and click the “submit” button to send the query back to the system. The user input data is stored in a designated field and sent to the backend, which then triggers the kql_executor skill to execute the query on behalf of the user.
@@ 7.2. Access Control
Access control for the bot is managed through Azure Active Directory (AAD) groups (Microsoft, 2024k). Although the web API endpoint is publicly accessible (necessary for communication with Teams, direct link channels, and potential future integrations), the bot itself handles the authentication process. The bot uses OAuth 2.0 (Microsoft, 2024l), with Azure Active Directory as the identity provider. This setup involves registering the bot with Azure, allowing it to function as an application that requests access to user data.
@@ 7.3. Telemetry
The copilot emits telemetry data to a Log Analytics Workspace, making this data accessible to Azure resource owners. Table 2 shows the types of data that are logged.
@@ 8. Evaluation
The evaluation system in DECO supports the continuous development of the framework to enhance chat quality. The evaluation is divided into two categories: online and offline evaluation.
@@ 8.1. Online Evaluation
The online evaluation assesses the direct interactions between the chatbot and users, producing performance indicators for actual conversations. In addition to user feedback ratings, we employ the following metrics: (1) relevance of the generated answer, scored from 1 to 3, (2) relevance of the retrieved documents, scored from 1 to 3, and (3) groundedness of the answer based on the retrieved documents, scored from 0 to 1. Each metric is evaluated using a specific prompt and an LLM to generate a numeric score. Additionally, an AML endpoint has been deployed to asynchronously collect supplementary telemetry data alongside the primary backend flow. We have categorized the results into five categories to clarify these scores for users (see Table 3).
@@
The online evaluation offers valuable metrics to compare the quality across various deployments for a range of Microsoft products. It also enables the measurement of document retrieval quality, which facilitates communication with the owning teams about areas where documentation can be further improved or is currently lacking.
8.2. Offline Evaluation
To support ongoing development of the chatbot and evaluate changes such as prompt modifications and updates to retrieval algorithms, we developed an evaluation framework that allows testing in hypothetical scenarios without disrupting the service. The offline evaluation focuses on assessing the quality of chatbot responses after code changes. We observed that:
Observation: LLM evaluation often differs significantly from human judgments, especially those made by domain experts, in terms of relevance and usefulness.
Using a large sample of user ratings, we explored the feasibility of employing an LLM to rate chatbot responses on a scale from 1 to 5 stars, aiming at simulating human judgment. When comparing LLM-generated ratings with human ratings across 54 questions, we found a correlation of only 0.35 (see Figure 8), indicating substantial discrepancies between LLM and human evaluations. Manual reviews revealed that many responses deemed inaccurate by domain experts were perceived as plausible by non-experts due to subtle errors that only specialists could detect. This highlights the limitations of using LLMs as evaluators for assessing chatbot responses, as discussed in (Liu et al., 2023). As a result, we prioritize using question-answer pairs with user ratings as benchmarks for evaluating chatbot responses whenever feasible.
Evaluators. In this paper, we discuss four commonly used evaluators: the TSG Evaluator, the Incident Retrieval Evaluator, the Planner Evaluator, and the Similarity Evaluator.
Figure 8. Comparison of human feedback with LLM-predicted ratings
TSG Evaluator
The TSG Evaluator is designed to assess the accuracy of documentation chunks retrieved by the TSG retrieval process that are subsequently referenced in the chatbot’s responses, i.e., by the get_tsg skill. The first “golden” set of TSG documentation is compiled from real user interactions, focusing only on responses rated four or five stars based on the “reference list” at the end of the chatbot’s response, which includes clickable URLs to the cited documents as directed by the prompt. The second “golden” set is generated by the LLM, which creates synthetic questions for each documentation chunk, allowing us to identify the corresponding TSG documentation that answers those questions, similar to the self-supervised learning (Wang and Castro Fernandez, 2023).
The evaluation process involves the following steps:
  • Calculating precision and recall metrics by comparing $G$ and $Q^{*}$:
    (3) $\text{Coverage}=\frac{|G\cap Q|}{|G|}.$
Figure 9. Comparison of different retrieval algorithms.
Figure 9 compares various retrieval algorithms as discussed in Section 6. The label “FT” indicates the use of document filtering, “RF” represents query rephrasing, and “HB” denotes hybrid semantic search. We tested 65 questions, with 28 sourced from real user data rated above four stars and the remaining 37 generated by LLM based on frequently retrieved document chunks. The results show that hybrid search (“HB-*”) significantly improves recall compared to other methods. Document filtering effectively reduces context length without compromising recall. Although query rephrasing does not notably enhance retrieval accuracy for these specific examples, it proves extremely useful for questions related to Incident Management tickets or those that are poorly formatted by users. For incident-related questions, where the raw query is combined with the incident ticket summary, search results are often suboptimal. However, rephrasing—similar to the HyDE process—enables the generation of optimized search text based on the ticket summary, improving document retrieval. Since generating ticket summaries involves an LLM call, the rephrasing process can be parallelized, avoiding additional latency on the critical path.
(a) Context length
(b) Average recall
Figure 10. TSG evaluator results.
Figure 10 illustrates the comparison of context length and average recall across three different embedding models from Azure OpenAI (OpenAI, 2024): “text-embedding-3-large,” “text-embedding-3-small,” and “text-embedding-ada-002”, applied to the real user queries. The x-axis represents the configuration settings for the maximum number of documents retrieved. The results show that the “text-embedding-3-large” model, with similar context lengths, yields more accurate retrieval outcomes. We conclude that even within the current framework:
Observation: Configuration tuning, including the choice of search algorithm (e.g., hybrid vs. vector search), selection of search fields (e.g., title or content), and embedding models, significantly impacts retrieval performance.
Incident Retrieval Evaluator
The evaluation process involves the following steps:
  • Collecting 100 incidents referred to in real user queries.
  • Extracting the summary of each incident.
We compare these results against the baseline “Similar Incident API” developed by the IcM portal. As shown in Figure 11, our incident retrieval method significantly outperforms the baseline, retrieving more incidents with a “High” similarity score. Furthermore, our customized re-ranking algorithm (discussed in Section 6) further improves performance, achieving a “High” similarity score for 97 out of 100 incident pairs.
Planner Evaluator
  • Executing the planner module for each question.
  • Retrieving the list of skills invoked based on the planner’s output.
Figure 12 shows the evaluation results for different planner algorithms discussed in Section 6.1. Due to the inherent uncertainty in LLM calls, each setting includes five experimental runs, each with 18 user questions and verified required skill sets. We compare the use of few-shot vs. zero-shot examples and flat vs. hierarchical planners. We observe significant improvements when few-shot examples are used, particularly when incorporating a hierarchical structure for selecting the best skills.
Figure 12. Evaluation of different planner algorithms.
Similarity Evaluator
To maintain consistency in response quality following code modifications, we rerun the chatbot to generate new responses and benchmark them against historically high-rated answers. Specifically, we used LLM to assess the similarity between two answers to the same question, using a 1-5 star rating scale. Each newly generated response aims for a similarity rating of at least 4.0 when compared to prior highly-rated answers.
Evaluation Framework. With the described evaluators and additional tools, we built an efficient and flexible offline evaluation framework (see Figure 13). The framework connects to the telemetry database using managed identity, allowing for the download of extensive real user queries, bot outputs, and user feedback. Additionally, a curated set of highly user-rated question-answer pairs serves as the “golden answer” in the evaluation process. The framework features a runner that uses a configuration file to define which evaluators to include in the evaluation pipeline. Each evaluator, inheriting from a base class, implements three main functions:
Figure 13. Offline evaluation framework
Users can extend the framework by adding more evaluators using the provided templates. The evaluation pipeline is integrated into the Azure Pipeline and triggered by pull requests.
@@ 9. Related Work
For automating error triage of IcM tickets, AutoTSG (Shetty et al., 2022) constructs executable scripts using TSG documentation and leverages a combination of machine learning and program synthesis to identify relevant troubleshooting steps tailored to specific issues. Previous works have addressed related challenges, such as automatic debugging and diagnosis of SQL query execution (Zhou et al., 2024), Spark jobs (Lu et al., 2019, 2017), and general root cause analysis for incidents (Roy et al., 2024; Zhang et al., 2024c; Ahmed et al., 2023; Zhang et al., 2024a; Chen et al., 2023; Demarne et al., 2020). However, these efforts primarily focus on specific types of issues. AskBrain (Chen et al., 2020) employs language models to automatically categorize and assign incidents to appropriate teams, reducing mitigation time. In contrast, our work addresses a broader range of problems, with the copilot supporting general queries beyond incident management to deliver accurate solutions.
To enhance retrieval-augmented generation (RAG) accuracy, TILDE (Zhuang and Zuccon, 2021b, a) and related works (Nogueira et al., 2019) propose re-ranking documents using fine-tuned likelihood or embedding models, as well as pretrained relevancy models (Nogueira et al., 2020). ReSP (Jiang et al., 2024) iteratively retrieves documents with a “Reasoner” to determine if the current set of documents is sufficient, although this method introduces significant latency. Similarly, (Ma et al., 2024) fine-tunes the LLaMA model as both a dense retriever (RepLLaMA) and a pointwise reranker (RankLLaMA) in a multi-stage text retrieval pipeline, achieving improved retrieval performance.

### v2

@@ 6.4. Code Retrieval
  • Search Field: A list specifying the scope of the search, which can include one or more of the following values:
@@
      title”: Searches the codebase by title, useful for locating where specific elements like methods or classes are defined.
@@
      description”: Searches the codebase by the LLM generated description, applicable when user asks for explanation of the code.
@@
      content”: Searches the codebase by content, applicable when seeking definitions of specific methods or classes.
    4. (4)
      reference”: Searches the codebase by reference, targeting locations where specific methods or classes are referenced.
@@ 7. Frontend
The frontend of DECO is built around the Azure Bot, which leverages the Azure Bot Framework V4 and is deployed on an Azure Web Application (see Figure 7). This setup coordinates API calls, manages messaging (interacting with the backend REST API), retains state in cloud storage, and supports user interface enhancements through frontend plugins such as Adaptive Cards (Microsoft, 2024a), allowing users to modify generated Kusto queries. Authentication is handled via Microsoft Entra ID (see Figure 7). The web application connects users to the backend Azure Bot through direct line web channels, which communicate with the Azure Bot API. This web channel seamlessly integrates with the IcM portal, pre-populating inputs like ticket IDs for backend processing. Additionally, the bot can be accessed via Microsoft Teams through a published manifest. The web application also transmits system telemetry to a reporting dashboard for monitoring. Detailed interactions between the components are shown in Figure 7.
@@ 7.1. Plugin
The frontend “Plugins” support various types of user interfaces required by certain skills. For example, for the kql.generator skill, the backend output includes plugin data that triggers a frontend plugin, which displays a special widget (an Adaptive Card) for the user. This widget allows users to select Kusto clusters from a dropdown menu, edit the query in the text box, use checkboxes to choose which queries to execute, and click the “submit” button to send the query back to the system. The user input data is stored in a designated field and sent to the backend, which then triggers the kql.executor skill to execute the query on behalf of the user.
@@ 7.2. Access Control
Access control for the bot is managed through Azure Active Directory (AAD) groups (Microsoft, 2024h). Although the web API endpoint is publicly accessible (necessary for communication with Teams, direct link channels, and potential future integrations), the bot itself handles the authentication process. The bot uses OAuth 2.0 (Microsoft, 2024i), with Azure Active Directory as the identity provider. This setup involves registering the bot with Azure, allowing it to function as an application that requests access to user data.
@@ 7.3. Telemetry
The copilot emits telemetry data to a cloud storage (i.g., Log Analytics Workspace), making this data accessible to copilot owners. Table 1 shows the types of data that are logged.
@@ 8. Evaluation
The evaluation system in DECO is designed to support the continuous improvement of the framework by (1) helping to triage the reasoning behind low-quality responses and (2) evaluating the performance of individual components to guide further development and optimization. The evaluation process is divided into two categories: offline evaluation and online evaluation.
8.1. Offline Evaluation
To support ongoing development of the copilot and evaluate changes such as prompt modifications and updates to retrieval algorithms or skill selection mechanisms, we developed an evaluation framework that allows testing in hypothetical scenarios without disrupting the service. The offline evaluation focuses on assessing the quality of copilot responses after code changes.
We observed that LLM evaluation often differs significantly from human judgments, especially those made by domain experts, in terms of relevance and usefulness. Using a sample of user ratings, we explored the feasibility of employing an LLM to rate copilot responses on a scale from 1 to 5 stars, aiming at simulating human judgment. When comparing LLM-generated ratings with human ratings across 54 questions, we found a correlation of only 0.35 (see Figure 9), indicating substantial discrepancies between LLM and human evaluations. Manual reviews revealed that many responses deemed inaccurate by domain experts were perceived as plausible by non-experts due to subtle errors that only specialists could detect. This highlights the limitations of using LLMs as evaluators for assessing copilot responses, as discussed in (Liu et al., 2023). As a result, we prioritize using question-answer pairs from real user conversations with user ratings as benchmarks for evaluating copilot responses whenever feasible.
Evaluation Framework. We built an flexible offline evaluation framework (see Figure 8) for all the evaluation/test suites. The framework connects to the telemetry database using managed identity, allowing for the download of extensive real user queries, bot outputs, and user feedback. A curated set of highly user-rated question-answer pairs serves as the “golden answer” in the evaluation process. The framework features a runner that uses a configuration file to define which evaluators to include in the evaluation pipeline. Each evaluator, inheriting from a base class, implements three main functions:
Figure 8. Offline evaluation framework
Users can extend the framework by adding more evaluators using the provided templates. The evaluation pipeline is integrated into the Azure Pipeline and triggered by pull requests (PRs).
Evaluators. In this paper, we discuss four commonly used evaluators: the Planner Evaluator, the TSG Retrieval Evaluator, the Incident Retrieval Evaluator, and the Similarity Evaluator.
Figure 9. Comparison of human feedback with LLM-predicted ratings
Planner Evaluator
  • Executing the planner module for each question.
  • Retrieving the list of skills invoked by the planner.
Whenever new skills/agents are added, we conduct experiments with this evaluator to ensure that the current planner can accurately select the skill when appropriate. If not, we fine-tune the skill’s description, the planner’s prompt, or the few-shot examples, to reach the optimal performance.
Figure 10 shows the evaluation results for different planner algorithms discussed in Section 5. Due to the inherent uncertainty in LLM calls, each setting includes five experimental runs, each with 18 representative user questions and verified required skill sets. We compare the use of few-shot vs. zero-shot examples and flat vs. agentic planners. We observe significant improvements when few-shot examples are used, particularly when incorporating a agentic structure for selecting the best skills with good consistency. In production, we also observed this behavior consistently.
Figure 10. Evaluation of different planner algorithms.
TSG Evaluator
The TSG Evaluator is designed to assess the accuracy of documentation chunks retrieved by the TSG retrieval process that are subsequently referenced in the copilot’s responses, i.e., by the get_tsg skill. The first “golden” set of TSG documentation is compiled from real user interactions, focusing only on responses rated four or five stars based on the “reference list” at the end of the copilot’s response, which includes clickable URLs to the cited documents as directed by the prompt. The second “golden” set is generated by the LLM, which creates synthetic questions for each documentation chunk, allowing us to identify the corresponding TSG documentation that answers those questions, similar to the self-supervised learning (Wang and Castro Fernandez, 2023).
The evaluation process involves the following steps:
  • Calculating precision and recall metrics by comparing $G$ and $Q^{*}$:
    (3) $\text{Coverage}=\frac{|G\cap Q|}{|G|}.$
Figure 11 compares various retrieval algorithms as discussed in Section 6. The label “FT” indicates the use of document filtering, “RF” represents query rephrasing, and “HB” denotes hybrid semantic search. We tested 65 questions where the developer team has access to, with 28 sourced from real user data rated above four stars and the remaining 37 generated by LLM based on frequently retrieved document chunks. The results show that hybrid search (“HB-*”) significantly improves recall compared to other methods. Document filtering effectively reduces context length without compromising recall. Although query rephrasing does not notably enhance retrieval accuracy for these specific examples, it proves extremely useful for questions related to Incident Management tickets or those that are poorly formatted by users. For incident-related questions, where the raw query is combined with the incident ticket summary, search results are often suboptimal. However, rephrasing—similar to the HyDE process—enables the generation of optimized search text based on the ticket summary, improving document retrieval. Since generating ticket summaries involves an LLM call, the rephrasing process can be parallelized, avoiding additional latency on the critical path.
Figure 11. Comparison of different retrieval algorithms.
Figure 12 illustrates the comparison of context length and average recall across three most recent embedding models from Azure OpenAI (OpenAI, 2024): “text-embedding-3-large,” “text-embedding-3-small,” and “text-embedding-ada-002”, applied to the real user queries. The x-axis represents the configuration settings for the maximum number of documents retrieved. The results show that the “text-embedding-3-large” model, with similar context lengths, yields more accurate retrieval outcomes.
(a) Context length
(b) Average recall
Figure 12. TSG evaluator results.
We conclude that even within the current framework, configuration tuning, including the choice of search algorithm (e.g., hybrid vs. vector search), selection of search fields (e.g., title or content), and embedding models, significantly impacts retrieval performance.
Incident Retrieval Evaluator
The evaluation process involves the following steps:
  • Collecting incident IDs referred to in real user queries.
  • Extracting the summary of each incident.
We executed the evaluator against 100 sampled user questions. We compare these results against the baseline “Similar Incident API” developed by the IcM portal. As shown in Figure 13, our incident retrieval method significantly outperforms the baseline, retrieving more incidents with a “High” similarity score. Furthermore, our customized re-ranking algorithm (discussed in Section 6) further improves performance, achieving a “High” similarity score for 97 out of 100 incident pairs.
Similarity Evaluator
To maintain consistency in response quality following code modifications, we rerun the copilot to generate new responses and benchmark them against historically high-rated answers. Specifically, we used LLM to assess the similarity between two answers to the same question, using a 1-5 star rating scale. Each newly generated response aims for a similarity rating of at least 4.0 when compared to prior highly-rated answers.
@@ 8.2. Online Evaluation
The online evaluation assesses the direct interactions between the copilot and users, producing performance indicators for actual conversations. In addition to user feedback ratings, we employ the following metrics: (1) relevance of the generated answer, scored from 1 to 3, (2) relevance of the retrieved documents, scored from 1 to 3, and (3) groundedness of the answer based on the retrieved documents, scored from 0 to 1. Each metric is evaluated using a specific prompt and an LLM to generate a numeric score. Additionally, an AML endpoint has been deployed to asynchronously collect supplementary telemetry data alongside the primary backend flow. We have categorized the results into five categories to clarify these scores for users (see Table 2).
@@
For telemetry data collected through online evaluation (enabled by a subset of teams), we analyzed over 3,000 messages and categorized the responses as shown in Figure 14 from 2025-01-01 to 2025-02-28. The copilot averages 2.2 interaction rounds per session, with 6 rounds at the $95^{\text{th}}$ percentile and 11 rounds at the $99^{\text{th}}$ percentile, indicating that users often engage in multi-turn conversations. The most common issue identified was Documentation Issues (23%), while hallucination cases (i.e., Grounding Issues) accounted for 0.1%. Among conversations flagged as having documentation issues, the root causes can be attributed to three potential factors:
  1. (1)
    Missing documentation – The relevant document does not exist in the knowledge base.
  2. (2)
    Retrieval skills not being invoked – The planner fails to trigger the appropriate retrieval skills.
  3. (3)
    Irrelevant retrieved documents – The retrieved documents do not align with the user query.
Among responses identified as documentation issues, 68% of these responses state upfront that “no relevant information was found”, while 27% provide a reference list of documents for user verification. This suggests that the LLM effectively detects irrelevant retrieved documents and appropriately disregards them.
Figure 14. Online evaluation category distribution
Combining user feedback, among the responses where users gave a low rating ($\leq$ 1 star), 44% were identified with documentation issues. Among these, 93% explicitly indicated that no relevant information was found. Other low-rating responses where users provided text feedback highlight potential areas for copilot improvement, such as missing functionalities (e.g., inability to execute code) and off-topic responses. These issues may stem from inaccuracies in documentation retrieval or ambiguous user input.
@@ 9. Related Work
For automating error triage of IcM tickets, AutoTSG (Shetty et al., 2022) constructs executable scripts using TSG documentation, machine learning, and program synthesis to identify relevant troubleshooting steps. Prior works have also focused on automatic debugging and diagnosis of SQL query execution (Zhou et al., 2024), Spark jobs (Lu et al., 2019, 2017), and general root cause analysis for incidents (Roy et al., 2024; Zhang et al., 2024c; Ahmed et al., 2023; Zhang et al., 2024a; Chen et al., 2023; Demarne et al., 2020). AskBrain (Chen et al., 2020) applies language models to automatically categorize and assign incidents to appropriate teams, reducing mitigation time. However, these efforts primarily target specific categories of issues.
To improve retrieval-augmented generation (RAG) accuracy, TILDE (Zhuang and Zuccon, 2021b, a) and related works (Nogueira et al., 2019) propose re-ranking documents using fine-tuned likelihood or embedding models, as well as pretrained relevancy models (Nogueira et al., 2020). ReSP (Jiang et al., 2024) iteratively retrieves documents using a “Reasoner” to assess the current retrieved documents, although this introduces significant latency. (Ma et al., 2024) fine-tunes the LLaMA model as both a dense retriever and a re-ranker in a multi-stage text retrieval pipeline, achieving improved retrieval performance. Retrieval-augmented methods have also been applied to code repositories for code completion tasks, such as RepoCoder (Zhang et al., 2023) and GraphCoder (Liu et al., 2024). However, these systems typically focus on generating code completions from partially written code, rather than handling natural language queries.
Recent research has also explored agentic frameworks to manage complex workflows involving multiple tools and skills. For example, systems such as AutoGen (Wu et al., 2023) and Letta (Packer et al., 2023) introduce multi-agent coordination patterns for task planning and execution, primarily in research or prototype environments. Gorilla (Patil et al., 2023) focuses on large-scale API selection by prompting language models to generate function calls from extensive catalogs, demonstrating the potential of LLMs for dynamic skill invocation. LlamaIndex (Liu and contributors, 2022) provides a general-purpose framework for connecting LLMs with external knowledge sources. In contrast, DECO adopts a lightweight agentic framework optimized for production-scale deployments, decoupling backend and frontend interactions. This design improves scalability and adaptability, particularly as the number of available skills grows in enterprise environments.